The total cost in thousands of shillings of producing q units of a given product is given by the function 1000+ 2q^2-12q^2.a) total fixed cost



b).The output level that will minimize the marginal cost.



c).The marginal cost when the level of output is 5000 units

1
Expert's answer
2022-01-24T10:54:01-0500

a) total fixed cost

TFC = Sh 1,000,000


b).The output level that will minimize the marginal cost.

If TC = 1000+2q2 -12q

MC = 4q - 12

At minimum MC: 4q - 12 = 0

4q = 12

q =3

Output level = 3000 units


c).The marginal cost when the level of output is 5000 units

MC = 4q - 12

at 5000 units: MC==> 4(5000) - 12000

MC = 20,000 -12,000


MC =sh 8,000